solve the following equation
-4/13x=-20

Answer by jim_thompson5910(35256)

-4/13x=-20

-4x=-20*13

-4x=-260

x=-260/(-4)

x = 65


So the solution is x = 65


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim